OpenStudy (anonymous):

lim sin(x-1)/x^2 +x-2 x->1

OpenStudy (anonymous):

factor the denom. : \[\frac{ \sin(x-1) }{ (x-1)(x+2) }\] \[\lim_{x-1 \rightarrow 0}\frac{ \sin(x-1) }{ (x-1) }*\lim_{x \rightarrow1} \frac{ 1 }{ (x+2) }\]